Dumbgirl2:

The company discovers that it costs $29 to produce 2 widgets, $115 to produce 4 widgets, and $757 to produce 10 widgets. Complete the function that represents the cost, c (x), to produce x widgets. C (x)=?

Ultrilliam:

Looks like this question has been asked before! questioncove.com/updates/55573d0ee4b0b98f70b37242 Does this help you at all?

Dumbgirl2:

I know the steps to the equation. When I did the equation, i got 8x^2-5x, but when I typed my answer in, it said it was wrong, and the correct answer was 8x^2-5x+7. I'm not sure how to get 7 into the equation

Ultrilliam:

Hero should be able to help you out with that, *he is looking at the question, but it may be a little bit before he responds, hence me posting this ¯\_(ツ)_/¯*

Dumbgirl2:

Ok, thank you

Hero:

@Dumbgirl2, I have a solution for this, but I curious, how did you know the function was quadratic?

Dumbgirl2:

Because I'm helping my boyfriend with his math, and there is a tutorial you do first, and it tells you at the top what type of function it is

Hero:

"It tells you at the top what kind of function it is". May I see what it is you're looking at? Can you screenshot it and upload it here if you don't mind?

Ultrilliam:

And easy way to do that would be by hitting the print screen key on your keyboard while looking at it, then simply hitting ctrl+v here while typing a reply, and the image will be auto-uploaded ^_^

Hero:

Test

1 attachment
Hero:

Holy wow. That's awesome.

Hero:

Awesome feature

Ultrilliam:

I figured that would simplify things :P

Hero:

You should do a tutorial for these cool features.

Ultrilliam:

Perhaps I should. Perhaps I should.

Dumbgirl2:

1 attachment
Ultrilliam:

Uhm, It's a wee bit low resolution o-o

Hero:

Thanks @Dumbgirl2 but I was hoping you could post a screenshot of the actual problem itself.

Dumbgirl2:

I already finished the tutorial, but when I started doing practice questions, they made no sense to me, and I even took notes on how to do it, but those don't help. And I can go back and try to get the same one. Sometimes the questions will be different

Hero:

That's okay, it can be different. I just wanted to observe the structure of the question itself as given from the original source.

Dumbgirl2:

Ok, it may be a while until I get question like that again

Hero:

I'm a very patient person.

Dumbgirl2:

Ok, thank you

Dumbgirl2:

1 attachment
Hero:

I see. So it lets you know that you are doing questions on applications of quadratic functions. So that's good because to be honest, I couldn't discern from the question alone what kind of function we're dealing with so that's good.

Hero:

So the way to go about solving this is to use the fact that a quadratic function, in this case, has the form: C(x) = ax^2 + bx + c Now the give you basically three points for (x,C(x)) which are (2,29), (4, 115) and (10,757) So you take those points and plug them in to the quadratic function like so: 29 = a(2)^2 + b(2) + c 115 = a(4)^2 + b(4) + c 757 = a(10^2 + b(10) + c And all that simplifies to: 29 = 4a + 2b + c 115 = 16a + 4b + c 757 = 100a + 10b + c So what we have is a system of equations in three variables to solve.

Hero:

To solve for a, b, and c. Solving for that should give you a = 8, b = -5, and c = 7

Hero:

Hope that answers your question.

Dumbgirl2:

How do you get 7

Dumbgirl2:

A and B I can usually get, but lately I haven't been getting the answer to C, and I don't know how to get it

Hero:

Okay, so assuming the values for a and b are correct, we should be able to simply take one of the equations above, input the values for a and b, then solve for c. Let's see what happens: 29 = 4a + 2b + c 29 = 4(8) + 2(-5) + c 29 = 32 - 10 + c 29 = 22 + c c = 29 - 22 c = 7

Dumbgirl2:

Oh my gosh thank you, this all makes sense now

Hero:

Koolaid

Hero:

There's another way to get c

Hero:

Well, basically, it would involve either solving the system fully or converting the system to matrices and solving it that way. But what I've shown above is definitely the easiest way.

Dumbgirl2:

And the quickest probably

Dumbgirl2:

Well at least the quickest for me

Hero:

Oh yes, definitely. If you already know a and b, then apply the strategy above, and finding c should be a piece of cake.

Dumbgirl2:

It was, thank you

Hero:

Cool. If you have any other questions, feel free to post them after closing this post. Otherwise, I hope you enjoy the rest of your evening. And good luck helping your friend.

Hero:

flat,800x800,075,f.u3.jpg

Hero:

@Dumbgirl2

Dumbgirl2:

1 attachment
Hero:

Do you know how to post this as a new question?

Dumbgirl2:

Not unless I go to ask a question and type the question out

Hero:

Yes you don't have to type the question out completely. Just write New Question

Hero:

Then post

Dumbgirl2:

Ok